Find the total work done in moving a particle in a force field given by 𝐹 = 3π‘₯𝑦𝑖 βˆ’ 5𝑧𝑗 + 10π‘₯π‘˜ along the curve x = 𝑑2 + 1, 𝑦 = 2𝑑2 , 𝑧 = 𝑑3 π‘“π‘Ÿπ‘œπ‘š 𝑑 = 1 π‘‘π‘œ 𝑑 = 2.


1
Expert's answer
2021-05-28T10:23:29-0400

"\begin{array}{l} {\overline{F}.\overline{dr}=(3xyi-52j+10x\overline{k})(dxi+dyi+dzk)} \\ {\, \, \, \, \, \, \, \, \, \, =3xydx-52dy+10xdz} \\ {\, \, \, \, \, \, \, \, \, \, \, =\, 3(t^{2} +1)(2t^{2} )(2tdt)-5(t^{3} )(4tdt)+10(t^{2} +1)(3t^{2} dt)} \\ {\, \, \, \, \, \, \, \, \, \, \, =(12t^{5} +12t^{3} -20t^{4} +30t^{4} +30t^{2} )dt} \\ {\, \, \, \, \, \, \, \, \, \, =(12t^{5} +10t^{4} +12t^{3} +30t^{2} )dt} \end{array}" "\begin{array}{l} {Work\, \, Done=\int _{C}f.\overline{dr} =\int _{1}^{2}(12t^{5} +10t^{4} +12t^{3} +30t^{2} )dt } \\ {\, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, \, =\, \left[\frac{12t^{6} }{5} +\frac{10t^{5} }{5} +\frac{12t^{4} }{4} +\frac{30t^{3} }{3} \right]_{1}^{2} } \\ {\qquad \, \, \, \, \, \, \, \, \, =\left[2t^{6} +2t^{5} +3t^{4} +10^{3} \right]_{1}^{2} } \\ {\qquad \, \, \, \, \, \, \, \, \, \, =\, \left[2(2)^{6} +2(2)^{5} +3(2)^{4} +10(2)^{3} \right]-\left[2+2+3+10\right]} \end{array}" "= 303"